### v1

# Myers–Steenrod theorems for metric and singular Riemannian foliations

Abstract.
We prove that the group of isometries preserving a metric foliation on a closed Alexandrov space $X$, or a singular Riemannian foliation on a manifold $M$ is a closed subgroup of the isometry group of $X$ in the case of a metric foliation, or of the isometry group of $M$ for the case of a singular Riemannian foliation. We obtain a sharp upper bound for the dimension of these subgroups and show that, when equality holds, the foliations that realize this upper bound are induced by fiber bundles whose fibers are round spheres or projective spaces. Moreover, singular Riemannian foliations that realize the upper bound are induced by smooth fiber bundles whose fibers are round spheres or projective spaces.
Key words and phrases:
Alexandrov space, submetry, isometry group, singular Riemannian foliation, Lie group
2020 Mathematics Subject Classification:
53C12,53C20,53C21,53C23,53C24,51K10
1 Supported in part by UNAM-DGAPA Postdoctoral fellowship of the Institute of Mathematics, and by the DFG (grant CO 2359/1-1, Priority Programme SPP2026 “Geometry at Infinity”), and by a UKRI Future Leaders Fellowship [grant number MR/W01176X/1; PI: J Harvey].
2 Supported in part by the DFG (grant GA 2050 2-1, Priority Programme SPP2026 “Geometry at Infinity”).
1. Main results
The Myers–Steenrod theorem [37] states that the isometry group of a Riemannian $n$-manifold is a Lie group whose dimension is at most $n(n+1)/2$. When the manifold is compact, its isometry group must also be compact, as established by van Dantzig and van der Waerden [13]. The application of the theory of compact transformation groups in Riemannian geometry [1, 7, 21, 28] is grounded on these two fundamental results, which also hold for other classes of metric spaces, such as Finsler manifolds [14], Alexandrov spaces [17, 18] or $\mathrm{RCD}$ spaces [24, 39].
Singular Riemannian foliations generalize both isometric compact Lie group actions and Riemannian submersions, which induce decompositions into embedded submanifolds of lower dimension, and represent a generalized notion of symmetry on Riemannian manifolds [3, 10, 12, 19, 36]. Not all singular Riemannian foliations stem from Lie group actions (see, for example, [38]). Nevertheless, certain isometries of a Riemannian manifold $M$ with a singular Riemannian foliation $\mathcal{F}$ may induce residual symmetry by preserving the foliation’s leaves. These foliated isometries generate isometries of the leaf space $M/\mathcal{F}$. Our first main result is an analogue of the Myers–Steenrod theorem for the group of foliated isometries of a Riemannian manifold with a singular Riemannian foliation.
Theorem A.
Let $(M,\mathcal{F})$ be a singular Riemannian foliation with closed leaves on a complete connected Riemannian manifold. Then the following assertions hold:
  1. (i)
    The group $\operatorname{Isom}(M,\mathcal{F})$ of foliated isometries is a Lie group and is compact if $M$ is compact.
  2. (ii)
    If $M$ has dimension $n\geq 1$ and $\mathcal{F}$ has codimension $0\leq k\leq n$, then
    (a) $\dim(\operatorname{Isom}(M,\mathcal{F}))\leq\frac{k(k+1)}{2}+\frac{(n-k)(n-k+1)}{2}.$
  3. (iii)
    If equality holds in inequality (a), then $M$ is foliated-diffeomorphic to a fiber bundle $F\rightarrow M\to B$, where $B$ is diffeomorphic to $\mathbb{R}^{k}$, $\mathbb{R}P^{k}$, or $S^{k}$, and the fibers are round spheres or real projective spaces. Moreover, when equipped with the leaf-projection metric induced by $M$, the base space $B$ is isometric to $k$-dimensional Euclidean space, hyperbolic space with constant negative sectional curvature, a round real projective space, or a round sphere.
The leaf space $M/\mathcal{F}$ of a singular Riemannian foliation $(M,\mathcal{F})$ with closed leaves has a natural distance function that makes $M/\mathcal{F}$ a locally compact length space. The curvature of $M/\mathcal{F}$ is locally bounded from below in the triangle comparison sense, which links the geometry of the leaf space with that of the manifold. This fact plays a central role in proving Theorem A. For an isometric action of a compact Lie group $G$ on a complete Riemannian manifold $M$ with sectional curvature uniformly bounded from below by $k\in\mathbb{R}$, the orbit space $M/G$ equipped with the orbital distance function is an Alexandrov space with curvature bounded below by $k$. Imposing further conditions, such as positive or non-negative sectional curvature on $M$, leads to significant constraints on the manifold’s topology and has been an active research topic in Riemannian geometry [21, 22, 40]. Note that the orbit projection map $\pi\colon M\to M/G$ is a proper submetry, i.e., for every $p\in M$, any closed metric ball $B(p,r)$ of radius $r$ centered at $p$ maps onto the metric ball $B(\pi(p),r)$ in $M/G$. Submetries, introduced by Berestovskii in [5] as a metric generalization of Riemannian submersions, have been the focus of systematic study in metric geometry [4, 25, 27, 29].
Motivated by the preceding considerations, we also investigate the foliated isometries of the metric foliations $\mathcal{F}$ whose leaves are the fibers of proper submetries $\pi\colon X\to Y$ between Alexandrov spaces. By the work of Lytchak [29], the leaves of such foliations must be closed. We consider the group $\operatorname{Isom}(X,\mathcal{F})$ of foliated isometries and obtain the following analogue of Theorem A.
Theorem B.
Let $\pi\colon X\to Y$ be a proper submetry between connected Alexandrov spaces and set =$\mathcal{F}=\{\pi^{-1}(y)\mid y\in Y\}$. Then the following assertions hold:
  1. (i)
    The group $\operatorname{Isom}(X,\mathcal{F})$ of foliated isometries of $(X,\mathcal{F})$ is a Lie group and is compact if $X$ is compact.
  2. (ii)
    If $X$ has dimension $n\geq 1$ and $Y$ has dimension $0\leqslant k\leqslant n$, then
    (b) $\dim(\operatorname{Isom}(X,\mathcal{F}))\leq\frac{k(k+1)}{2}+\frac{(n-k)(n-k+1)}{2}.$
  3. (iii)
    If equality holds in inequality (b), then $X$ is foliated-homeomorphic to a fiber bundle $F\rightarrow X\to Y$, where $Y$ is homeomorphic to $\mathbb{R}^{k}$, $\mathbb{R}P^{k}$, or $S^{k}$, and the fibers are isometric to a finite disjoint union og round spheres or projective spaces. Moreover, when equipped with the leaf-projection metric induced by $X$, the base space $Y$ is isometric to $k$-dimensional Euclidean space, hyperbolic space with constant negative sectional curvature, a round real projective space, or a round sphere.
The upper bound for the dimension of the group of foliated isometries in Theorems A and B consists of two summands. The first summand, $k(k+1)/2$, bounds the dimension of the group of foliated isometries $f$ that do not leave all leaves invariant, i.e., there exists some leaf $L\in\mathcal{F}$ with $f(L)\not\subset L$. These are the isometries that descend to non-trivial isometries of the leaf space. The second summand, $(n-k)(n-k+1)/2$, bounds the dimension of the group of foliated isometries $h$ that leave the leaves invariant, i.e., $h(L)\subset L$ for any leaf $L\in\mathcal{F}$.
Applying Theorem A to the trivial foliations =$\mathcal{F}=\{M\}$ or =$\mathcal{F}=\{{p}\mid p\in M\}$ consisting, respectively, of a single leaf or leaves that contain only one point, yields the classical upper bound $n(n+1)/2$ in the Riemannian Myers–Steenrod theorem. For Alexandrov spaces, Theorem B applied to the trivial submetries $\mathrm{id}\colon X\to X$ and $\pi\colon X\to\{\mathrm{pt}\}$ yields the bound $n(n+1)/2$ in the Myers–Steenrod theorem for Alexandrov spaces in [18].
As in the Riemannian, Alexandrov, and $\mathrm{RCD}$ Myers–Steenrod theorems, the upper bound in Theorems A and B is sharp. Consider the Riemannian product $(S^{n-k}\times S^{k},g_{1}\times g_{2})$ of two round spheres $(S^{k},g_{1})$, $(S^{n-k},g_{2})$ of different radii, with $n>2k$, and the Riemannian foliation $\mathcal{F}$ whose leaves are $S^{n-k}$. Then $\operatorname{Isom}(M,g)=\operatorname{Isom}(S^{k},g_{1})\times\operatorname{Isom}(S^{n-k},g_{2})$, and thus $\operatorname{Isom}(M,\mathcal{F})=\operatorname{Isom}(M)$. This realizes the upper bound in Theorem A (a) and Theorem B (b).
When equality holds in inequality (a), $M$ is a fiber bundle $F\to M\to B$ and we may assume that the leaves $F$ carry a metric of constant sectional curvature equal to one. The main theorems in [11] and [15] imply that, when $B$ is not diffeomorphic to $\mathbb{R}^{k}$, the bundle has a linear structure group. When $B$ is diffeomorphic to $\mathbb{R}^{k}$, the manifold $M$ is diffeomorphic to $\mathbb{R}P^{n-k}\times\mathbb{R}^{k}$ or $S^{n-k}\times\mathbb{R}^{k}$, since $\mathbb{R}^{k}$ is contractible. This last conclusion also holds for the case of Theorem B (iii) in the case when $Y$ is homeomorphic to $\mathbb{R}^{k}$.
The map $\operatorname{Isom}(M,\mathcal{F})\to\operatorname{Isom}(M/\mathcal{F})$ is generally not surjective. However, if $M$ is an $n$-dimensional Euclidean vector space with a foliation $\mathcal{F}$ induced by a linear isometric action by a compact Lie group $G$, any isometry in $\operatorname{Isom}(M/G)_{o}$ lifts to a foliated isometry in $\operatorname{Isom}(M,G)$ [33]. Here, $\operatorname{Isom}(M/G)_{o}$ denotes the connected component containing the identity. The proof of Theorem B yields an upper bound on the dimension of the orbits of $G$ in terms of the dimension of $M/G$. Specifically, if $M/G$ has dimension $0<k<n$, then an orbit has dimension at most $(n-k)(n-k+1)/2$.
Remark 1.1.
We point out that in the proof of Theorem A, we only use the transnormal properties of the closed foliation. Nonetheless, we state our results for singular Riemannian foliations, as there are no known examples of transnormal systems that are not smooth foliations, and it is conjectured that any transnormal system is a singular Riemannian foliation (see [40, Final Remarks]).
Remark 1.2.
In the proof of Theorem A (i), we heavily rely on the assumption that the leaf space is Hausdorff. To prove item Theorem A (ii), we require the leaf space to be an Alexandrov space. Both conditions are satisfied if the singular Riemannian foliation has closed leaves. However, this requirement can be relaxed by asking that the leaves of the foliation be globally equidistant instead of just locally equidistant.
Remark 1.3.
We do not know if the upper bound on the dimesion stated in Theorem A (ii) holds for singular Riemannian foliations without closed leaves. Even if such a bound exists, we cannot obtain a rigidity conclusion as in Theorem A (iii). This is illustrated by the irrational flow on the $2$-dimensional flat torus: the group of foliated isometries has dimension two, which is the dimension of the isometry group of the $2$-dimensional flat torus. In contrast, when we assume that the leaves are closed and the singular Riemannian foliation is not trivial, Theorem A (iii) implies that the foliation is given by a circle bundle over the circle.
Remark 1.4.
In þB (iii), there are examples of foliations with disconnected fibers. Namely, consider the product of a round $2$-sphere with a round $3$-sphere, and conisder the $3$-sphere as leaves. As remarked above, the group of foliated isometries has maximal dimension. Now consider the antipodal action of $\mathbb{Z}_{2}$ on $S^{2}$. With this, we get a fibration of $S^{2}\times S^{3}$ over $\mathbb{R}P^{2}$ whose fibers are two disjoint copies of a round $S^{3}$. The difference with þA is that we ask the leaves of a singular Riemannian foliation to be connected.
Our article is organized as follows. Section 2 presents basic material on singular Riemannian foliations, Alexandrov spaces, and submetries, as well as auxiliary results used in the proofs of our main theorems. Section 3 contains the proof of Theorem A. Finally, in Section 4, we prove Theorem B.
Acknowledgements.
We thank Alexander Lytchak and Marco Radeschi for helpful comments on a preliminary version of this article. We thank the organizers of the IV joint meeting of RSME and SMM, and the Universidad Politécnica de Valencia, for their hospitality while this manuscript was finished.
2. Preliminaries
In this section, we collect basic definitions and results on singular Riemannian foliations, Alexandrov spaces, and submetries we will use in the proof of our main theorems. We refer the reader to [2, 20, 8, 29, 27] for further basic results on theses subjects. We will assume all spaces to be connected, unless stated otherwise.
2.1. Singular Riemannian foliations
A singular Riemannian foliation $(M,\mathcal{F})$ on a complete Riemannian manifold $M$ is a partition of the manifold into a collection =$\mathcal{F}=\{L_{x}\mid x\in M\}$ of disjoint connected, complete, immersed submanifolds $L_{x}$, called leaves, satisfying the following conditions:
  1. (i)
    If $\gamma\colon[a,b]\to M$ is a geodesic perpendicular to the leaf $L_{\gamma(a)}$, then $\gamma$ is perpendicular to $L_{\gamma(t)}$ for all $t\in[a,b]$.
  2. (ii)
    For each point $x\in M$, there exists a smooth family of vector fields spanning $T_{x}L_{x}$, the tangent space to $L_{x}$ at $x$.
We call any leaf of maximal dimension a regular leaf; leaves that are not regular are called singular. Given any Riemannian manifold $M$, the foliation consisting of one leaf =$\mathcal{F}=\{M\}$ and the foliation where each leaf consists of just a point =$\mathcal{F}=\{\{x\}\mid x\in M\}$ are trivial examples of singular Riemannian foliations. Other examples of singular Riemannian foliations are given by the partition of a complete Riemannian manifold $M$ by the orbits of an isometric action of a compact Lie group. The partition of $M$ into the fibers of a Riemannian submersion $f\colon M\to N$ yields a further example of a singular Riemannian foliation. Note that there are infinitely many examples of singular Riemannian foliations which not given by group actions nor from Riemannian submersions [16, 38].
Let $(M,\mathcal{F})$ be a singular Riemannian foliation. Then we have a singular distribution $H\subset M$, called the horizontal distribution, given by setting $H_{x}=\nu_{x}L_{x}$, the normal tangent space to the leaf $L_{x}$ at $x$. The codimension of the foliation, denoted by $\operatorname{codim}(\mathcal{F})$, is the codimension of any regular leaf in $M$. If all the leaves of the foliation are closed submanifolds, we say that the foliation is closed. The leaf space of the foliation is the set of equivalence classes $M^{\ast}=M/\mathcal{F}$ equipped with the quotient topology. We have a natural projection map $\pi\colon M\to M^{\ast}$ which is continuous with respect to the quotient topology. Given a subset $A\subset M$, we let $A^{\ast}=\pi(A)$.
Let $(M,\mathcal{F})$ be a closed singular Riemannian foliation. Fix $x\in M$ and consider the normal tangent space $\nu_{x}L_{x}$ to the leaf $L_{x}\subset M$. Next, for $\varepsilon>0$ sufficiently small, define $\nu_{x}^{\varepsilon}L_{x}=\exp_{x}((\nu_{x}L_{x})\cap B_{\varepsilon}(0))$, where $B_{\varepsilon}(0)$ is the closed ball of radius $\varepsilon$ in $T_{x}M$. Set $S_{x}=\exp_{p}(\nu_{x}^{\varepsilon}L_{x})$. The intersection of the leaves of $\mathcal{F}$ with $S_{x}$ induces a foliation $\mathcal{F}|_{S_{x}}$ on $S_{x}$ whose leaves are the connected components of the intersection between the leaves of $\mathcal{F}$ and $S_{x}$. Although $\mathcal{F}|{S_{x}}$ may not be a singular Riemannian foliation with respect to the induced metric of $M$ on $S_{x}$ (the leaves of $\mathcal{F}|{S_{x}}$ may not be equidistant with respect to the induced metric), the pull-back foliation $\mathcal{F}^{x}=\exp_{x}^{\ast}(\mathcal{F}|_{S_{x}})$ is a singular Riemannian foliation on $\nu_{x}^{\varepsilon}L_{x}$ equipped with the Euclidean metric (see [35, Proposition 6.5]). The foliation $(\nu_{x}^{\varepsilon}L_{x},\mathcal{F}^{x})$ is called the infinitesimal foliation at $x$.
The infinitesimal foliation $(\nu_{x}^{\varepsilon}L_{x},\mathcal{F}^{x})$ is invariant under homotheties fixing the origin (see [35, Lemma 6.2]). Furthermore, the origin $\{0\}\subset\nu_{x}^{\varepsilon}L_{x}$ is a leaf of the infinitesimal foliation. Since the leaves of $\mathcal{F}^{x}$ are equidistant, the origin being a leaf implies that any leaf of $\mathcal{F}^{x}$ is at a constant distance from $\{0\}$. Therefore, each leaf of the infinitesimal foliation is contained in a round sphere centered at the origin. Hence, we may consider the infinitesimal foliation restricted to the unit normal sphere of $\nu_{x}L_{x}$, denoted by $S^{\bot}_{x}$, resulting in a foliated round sphere $(S^{\bot}_{x},\mathcal{F}^{x})$ with respect to the standard round metric of $S^{\bot}_{x}$. This foliation is also called the infinitesimal foliation. Henceforth, when referring to the ”infinitesimal foliation”, we will mean $(S^{\bot}_{x},\mathcal{F}^{x})$. Note that there is no loss of generality in doing so since $(\nu_{x}L_{x},\mathcal{F}^{x})$ is invariant under homothetic transformations and thus one may recover it from $(S^{\bot}_{x},\mathcal{F}^{x})$.
We denote by $\mathrm{O}(S^{\perp}_{x},\mathcal{F}^{x})$ the group of foliated isometries of the infinitesimal foliation, i.e., the isometries which preserve the infinitesimal foliation. For each loop $\gamma$ at $x$, the map $G_{\gamma}\colon\nu_{x}L\to\nu_{x}L$ given by $G_{\gamma}(v)=G(1,v)$ is a foliated linear isometry (see [34, Corollary 1.5]). Therefore, we have a group homomorphism $\rho\colon\Omega(L,x)\to\mathrm{O}(S^{\perp}_{x},\mathcal{F}^{x})$ from the loop space of $L_{x}$ at $x$ to the foliated isometries of the infinitesimal foliation by setting $\rho(\gamma)=G_{\gamma}$.
An isometry in $\mathrm{O}(S^{\perp}_{x},\mathcal{F}^{x})$ may map a leaf to a different leaf. By $\mathrm{O}(\mathcal{F}^{p})$ we denote the foliated isometries preserving the foliation, i.e., isometries $f\in\mathrm{O}(S^{\perp}_{x},\mathcal{F}^{x})$ such that for any leaf $\mathcal{L}$ of $(S^{\perp}_{x},\mathcal{F}^{x})$, we have $f(\mathcal{L})\subset\mathcal{L}$. The natural action of $\mathrm{O}(S^{\perp}_{p},\mathcal{F}^{x})$ on the quotient $S^{\perp}_{x}/\mathcal{F}^{x}$ has kernel $\mathrm{O}(\mathcal{F}^{x})$. Using the fact that if two loops $\gamma_{1}$ and $\gamma_{2}$ based at $x$ are homotopic, then $G_{\gamma_{1}}^{-1}\circ G_{\gamma_{2}}$ is in the kernel of the action of $\mathrm{O}(S^{\perp}_{x},\mathcal{F}^{x})$ on $S^{\perp}_{x}/\mathcal{F}^{x}$, one may show that there is a well-defined group homomorphism
$\rho\colon\pi_{1}(L,x)\to\mathrm{O}(S^{\perp}_{x},\mathcal{F}^{x})/\mathrm{O}(\mathcal{F}^{x}),$
given by $\rho[\gamma]=[G_{\gamma}]$ (see, for example, [10, Lemma 2.4 and Proposition 2.5]). We define the holonomy of the leaf $L$ as the image $\Gamma_{L}<\mathrm{O}(S^{\perp}_{x},\mathcal{F}^{x})/\mathrm{O}(\mathcal{F}^{p})$ of $\pi_{1}(L,p)$ under the homomorphism $\rho$. When we consider the holonomy of a leaf $L_{x}$ trough a point $x\in M$, we denote it by $\Gamma_{x}$. A regular leaf $L$ is a principal leaf if its holonomy group is trivial, and exceptional otherwise. The set $M_{\mathrm{prin}}\subset M$ consisting of the union of prinicpal leaves is an open, connected, and dense subset of $M$ (see [10, Theorem 3.12]).
2.2. Alexandrov spaces
An Alexandrov space $(X,d)$ with curvature bounded below by $k\in\mathbb{R}$ is a complete length space of finite Hausdorff dimension with curvature bounded below in the triangle comparison sense. Namely, for each $x\in X$, there is an open neighborhood $U\subset X$ of $x$, such that, for each geodesic triangle $\triangle$ contained in $U$, there exists a geodesic triangle $\widetilde{\triangle}$ in $M^{2}_{k}$, the $2$-dimensional model space of constant sectional curvature $k$, with edges having the same lengths as the edges of $\triangle$ satisfying the following condition: If $y$ is a vertex of $\triangle$, $\widetilde{y}$ is the corresponding vertex in $\widetilde{\triangle}$, and $w$ is any point in the opposing edge in $\triangle$ with corresponding point $\widetilde{w}$ in the opposing edge in $\widetilde{\triangle}$, then
$d(y,w)\geqslant d_{M^{2}_{k}}(\widetilde{y},\widetilde{w}).$
As is customary, we will abbreviate curvature bounded below by $k$ by writing $\operatorname{curv}\geqslant k$. A complete Riemannian manifold $M$ with a uniform lower bound for the sectional curvature is an example of an Alexandrov space.
Let $(X,d)$ be a metric space and fix three points $x,y,z\in X$. We define the comparison angle $\measuredangle(yxz)$ at $x$ as
$\measuredangle(y,x,z)=\arccos\left(\frac{d(x,y)^{2}+d(x,z)^{2}-d(y,z)^{2}}{2d(x,y)d(x,z)}\right).$
Now consider two continuous curves $c_{1}\colon[0,1]\to X$ and $c_{2}\colon[0,1]\to X$ with $c_{1}(0)=c_{2}(0)=x\in X$. We define the angle between $c_{1}$ and $c_{2}$ as
$\angle(c_{1},c_{2})=\lim_{s,t\to 0}\measuredangle(c_{1}(s),x,c_{2}(x)),$
provided the limit exists. When $X$ is an Alexandrov space and the curves $c_{1}$ and $c_{2}$ are geodesics, the angle $\angle(c_{1},c_{2})$ exists (see [8, Proposition 4.3.2]).
Given two geodesics $c_{1}\colon[0,1]\to X$ and $c_{2}\colon[0,1]\to X$ in an Alexandrov space $X$ with common start point $x\in X$, we say that $c_{1}$ is equivalent to $c_{2}$ if $\angle(c_{1},c_{2})=0$. Let $\widetilde{\Sigma}_{x}(X)$ the set of equivalence classes of geodesics starting at $x$. We define a metric on this set by setting the distance between two classes to be the angle formed between any two representatives of each class. The space of directions $\Sigma_{x}(X)$ of $X$ at $x$ is the metric completion of $\widetilde{\Sigma}_{x}(X)$. By Corollaries 7.10 and 7.11 in [9], the metric space $\Sigma_{x}(X)$ is an Alexandrov space of $\operatorname{curv}\geqslant 1$.
By Theorem 10.4.1 in [8], any Alexandrov space $\Sigma$ with $\operatorname{curv}\geqslant 1$ has $\operatorname{diam}(\Sigma)\leq\pi$. Let $(X,d)$ be a metric space with $\operatorname{diam}(X)\leq\pi$. The euclidean cone over $X$, denoted by $CX$, is the set $X\times[0,\infty)/(x,0)\sim(y,0)$ equipped with the metric
$d_{C}([x,t],[y,s])=\sqrt{t^{2}+s^{2}-2ts\cos(d(x,y))}.$
Given an Alexandrov space $X$ and a point $x\in X$, we define the tangent cone of $X$ at $x$ as $T_{x}X=C\Sigma_{x}(X)$. We denote the vertex of $T_{x}X$ by 00.
The following basic example illustrates a connection between the theories of Alexandrov spaces and of closed singular Riemannian foliations.
Example 2.1.
The leaf space $M^{\ast}$ of a singular Riemannian foliation $(M,\mathcal{F})$ on a complete Riemannian manifold with closed leaves inherits a complete metric $d^{\ast}$ from $M$, known as the leaf-projection metric. For $x^{\ast},y^{\ast}\in M^{\ast}$, the distance $d^{\ast}(x^{\ast},y^{\ast})$ is defined as
$d^{\ast}(x^{\ast},y^{\ast})=d(L_{x},L_{y}),$
the distance between the leaves $L_{x}$ and $L_{y}$ considered as subsets of $M$. Equipped with the metric $d^{\ast}$, the leaf space $M^{\ast}$ has curvature locally bounded below in the triangle comparison sense discussed above. Specifically, if $U\subset M$ is an open neighborhood with sectional curvature bounded below by $k_{U}\in\mathbb{R}$, then the projection $U^{\ast}\subset M^{\ast}$ has curvature bounded below by $k_{U}$ (see [31]). The Hausdorff dimension of $M^{*}$ is equal to the codimension of $\mathcal{F}$. Hence, if $M$ has sectional curvature uniformly bounded below by $k\in\mathbb{R}$, and $\mathcal{F}$ is closed, then $M^{*}$ is an Alexandrov space with $\operatorname{curv}\geqslant k$.
2.3. Submetries from proper metric spaces
Below, we collect several results from [29] on general submetries between Alexandrov spaces (see also [25] and [27]).
A map $p\colon X\to Y$ between two metric spaces is a submetry if, for any $\varepsilon>0$ and any $x\in X$, we have $p(B_{\varepsilon}(x))=B_{\varepsilon}(p(x))$. In other words, $p$ maps closed balls of radius $\varepsilon$ in $X$ onto closed balls of radius $\varepsilon$ in $Y$. Recall that a metric space $X$ is proper if every closed ball in $X$ is compact. Every proper metric space is complete.
Let $(X,d)$ be a metric space and let =$\mathcal{F}=\{L_{\alpha}\mid\alpha\in\Lambda\}$ be a partition of $X$. The foliation $\mathcal{F}$ is equidistant if, for any two distinct $L_{\alpha},L_{\beta}\in\mathcal{F}$, there exist $x_{\alpha}\in L_{\alpha}$ and $x_{\beta}\in L_{\beta}$ such that $d(x_{\alpha},x_{\beta})=d(L_{\alpha},L_{\beta})$. We will say that $\mathcal{F}$ is a metric foliation if it is an equidistant partition.
An element $L_{\alpha}\in\mathcal{F}$ of a metric foliation is a leaf. For any $x\in X$, we denote by $L_{x}$ the leaf of $\mathcal{F}$ containing $x$, and refer to $L_{x}$ as the leaf through $x$. The leaf space is the quotient space $X/\mathcal{F}$ whose elements are the leaves of the foliation. As for singular Riemannian foliations, we set $X^{*}=X/\mathcal{F}$. We define a metric on $X^{*}$ by letting $d(x^{*},y^{*})=d(L_{x},L_{y})$ for any $x^{*},y^{*}\in X^{*}$. The leaf-projection map $p\colon X\to X^{*}$ is then a submetry.
When $X$ is a proper metric space, the leaves of a metric foliation of $X$ are closed subsets of $X$. Moreover, if $p\colon X\to Y$ is a submetry between metric spaces, then the partition =$\mathcal{F}=\{p^{-1}(y)\mid y\in Y\}$ is a metric foliation of $X$ with closed leaves (see [29, p. 19]).
Lemma 2.2 ([29, Lemma 4.7]).
Let $p\colon X\to Y$ be a submetry. If $X$ is a proper metric space, then the following assertions are equivalent:
  1. (1)
    The map $p$ is proper.
  2. (2)
    The fibers of $p$ are compact.
  3. (3)
    There is a compact fiber.
Let $p\colon X\to Y$ be a submetry between two metric spaces. Two points $x_{1},x_{2}\in X$ are near (with respect to $p$) if $d(x_{1},x_{2})=d(p(x_{1}),p(x_{2}))$. The points $x_{1},x_{2}$ are near if and only if they realize the distance between the leaves $L_{x_{1}}$ and $L_{x_{2}}$ in $X$.
Consider a length space $X$, a metric space $Y$, and a submetry $p\colon X\to Y$. Let $\gamma$ be a geodesic in $Y$ through a point $y\in Y$. Given $x\in p^{-1}(y)$, there exists a geodesic $\tilde{\gamma}$ through $x$ that is mapped by $p$ isometrically onto $\gamma$. We call $\tilde{\gamma}$ the horizontal lift of $\gamma$. A geodesic in $X$ is horizontal if it is mapped by $p$ isometrically onto a geodesic in $Y$. A shortest path between two points $x_{1},x_{2}\in X$ is horizontal if and only if the points $x_{1},x_{2}$ are near (see [29, p. 21]).
Theorem 2.3 ([29, Theorem 7.2]).
Let $p\colon X\to Y$ be a submetry between metric spaces. Then the following assertions hold:
  1. (1)
    The components of each fiber $F_{y}=p^{-1}(y)$ are at distance at least $\varepsilon(y)$ from one another.
  2. (2)
    The intrinsic metric on each component of a fiber $F_{y}$ is locally Lipschitz-equivalent to the induced metric.
2.4. Submetries between Alexandrov spaces
When the metric space $X$ is an Alexandrov space, we may gain further insight into the local structure of submetries. To do so, we first recall several notions and results that will enable us to describe the space of directions of a family of leaves. We follow [29].
Let $(X,d)$ be a metric space. For each point $x\in X$, we let
$\operatorname{rad}_{x}=\sup_{y\in X}d(x,y),$
and define the radius of $X$, denoted by $\operatorname{rad}(X)$, by
$\mathrm{rad}(X)=\inf\{\mathrm{rad}_{x}\mid x\in X\}.$
Let $\Sigma$ be an $n$-dimensional Alexandrov space with curvature bounded below by $1$. Following [29, Definition 2.5], we will say that $\Sigma$ is Riemannian if it is isometric to the unit round $n$-sphere $S^{n}$. If $\operatorname{vol}(\Sigma)>(1-\rho)\operatorname{vol}(S^{n})$ for a fixed and sufficiently small positive real number $\rho=\rho(n)$, we refer to $\Sigma$ as extremely thick. If there exist $(n+2)$ points $v_{i}\in\Sigma$ with $d(v_{i},v_{j})>\pi/2$, we call $\Sigma$ thick. Additionally, we say $\Sigma$ is round if $\mathrm{rad}(\Sigma)>\pi/2$. By Grove and Petersen’s radius sphere theorem, every round Alexandrov space is homeomorphic to a sphere (see [23]). Given $A\subset\Sigma$, we let $\operatorname{Pol}(A)$ be the set of all $v\in\Sigma$ with $d(v,A)\geqslant\pi/2$, and refer to it as the polar set of $A$. Two points $v,w\in\Sigma$ are antipodal if $d(v,w)=\pi$. Toponogov’s comparison theorem implies that $\operatorname{Pol}(A)$ is a totally convex subset of $\Sigma$. Hence, $\operatorname{Pol}(A)$ is an Alexandrov space with $\operatorname{curv}\geqslant 1$.
Let $X$ and $Y$ be metric spaces and let $CX$ and $CY$ be their respective topological cones. A map $f\colon CX\to CY$ is homogeneous if
$f([t,x])=[t,f(x)]$
for all $t\in[0,\infty)$, and all $x\in X$. As mentioned in Section 2.2, if $\Sigma$ is an $n$-dimensional Alexandrov space with $\operatorname{curv}\geq 1$, then $C\Sigma$ may be endowed with a metric $d$ such that $(C\Sigma,d)$ is an $(n+1)$-dimensional Alexandrov space of non-negative curvature, the euclidean cone of $\Sigma$. We will refer to points in $C\Sigma$ as vectors. Given a vector $v=[t,u]$ in $(C\Sigma,d)$, we will refer to $t$ as the magnitude of $v$ and will set $|v|=t$. Let $p\colon C\Sigma\to C\mathrm{T}$ be a submetry between euclidean cones. A vector $h\in C\Sigma$ is horizontal (with respect to $p$) if $|p(h)|=|h|$.
Proposition 2.4 ([29, Proposition 6.4 and Lemma 6.5]).
Let $\Sigma$ and $\mathrm{T}$ be Alexandrov spaces with $\operatorname{curv}\geqslant 1$ and let $C\Sigma$, and $C\mathrm{T}$ be their respective euclidean cones. If $p\colon C\Sigma\to C\mathrm{T}$ is a homogeneous submetry, then the following assertions hold:
  1. (1)
    The preimage of the vertex $o\in C\mathrm{T}$ is a totally convex subcone $CV$ of $C\Sigma$ defined over a totally convex subset $V\subset\Sigma$.
  2. (2)
    The cone $CH\subset C\Sigma$ over $H=\operatorname{Pol}(V)\subset\Sigma$, the polar set of $V$, is the set of horizontal vectors.
  3. (3)
    Let $H=\operatorname{Pol}(V)$ as in item (2). Then $\operatorname{Pol}(H)=V$.
Let $V$ and $H$ be as in Proposition 2.4. Given $v\in V$ with $|v|=1$, consider the set
$H^{v}=\{h\in X\mid d(h,v)=\pi/2\}\subset\operatorname{Pol}(V)=H.$
The set $\widetilde{H}=\cap_{v\in V}H^{v}$ is the set of all points $h\in H$ that have distance equal to $\pi/2$ to any point in $V$. Since $H$ and $V$ are totally convex subsets of $\Sigma$, they are Alexandrov spaces of $\operatorname{curv}\geqslant 1$. Thus their spherical join $H*V$ is again an Alexandrov space with $\operatorname{curv}\geq 1$. Moreover, there exists a $1$-Lipschitz map $P_{H}\colon\Sigma\to H\ast V$ for which the following assertions are equivalent:
Proposition 2.5 ([29, Proposition 6.14 ]).
The following assertions are equivalent:
  1. (1)
    $P_{H}$ is surjective.
  2. (2)
    $H=\widetilde{H}$.
  3. (3)
    $P_{H}$ is a submetry.
If the submetry $P_{H}\colon\Sigma\to H\ast V$ satisfies any of the conditions in Proposition 2.5, we say that the set $H\subset\Sigma$ is almost spanning. Moreover, if $P_{H}$ is an isometry, we say that $H$ is spanning. We call a homogeneous submetry $p\colon C\Sigma\to C\mathrm{T}$ (almost) spanning if the horizontal set $H\subset\Sigma$ is (almost) spanning.
By considering the euclidean cones over the spaces $\Sigma$, $H\ast V$, $H$, and $\mathrm{T}$, and the maps between them discussed above, we may represent a submetry $p\colon C\Sigma\to C\mathrm{T}$ as a composition
[…]
where $\mathrm{pr}_{CH}$ is the projection onto the first factor of the metric product $CH\times CV$, and the other two maps are submetries.
Lemma 2.6 ([29, Lemma 6.15]).
Let $p\colon C\Sigma\to C\mathrm{T}$ be a homogeneous submetry and let $H\subset\Sigma$ be the set of horizontal vectors. If the restriction ${\left.\kern-1.2ptp\right|_{H}}\colon H\to\mathrm{T}$ is an isometry, then $H$ is almost-spanning.
A homogeneous submetry $p\colon C\Sigma\to C\mathrm{T}$ is regular if $p\colon H\to\mathrm{T}$ is an isometry.
Proposition 2.7 ([29, Lemma 6.16]).
Let $p\colon C\Sigma\to C\mathrm{T}$ be a homogeneous submetry. Then the following assertions hold:
  1. (1)
    If either $V$ or $H$ is a round space, then $p$ is spanning.
  2. (2)
    If $\mathrm{T}$ is a round space, then $p$ is regular and spanning.
We now recall the definition of differentiability for Lipschitz functions between Alexandrov spaces (cf. [29, Section 3]). Given a metric space $(X,d)$ and a real number $r>0$, we denote by $rX$ the metric space $(X,rd)$. Let $\{(X_{i},x_{i})\}_{i=1}^{\infty}$ and $(X,x)$ be pointed proper metric spaces. The sequence $\{(X_{i},x_{i})\}$ converges in the pointed Gromov–Hausdorff sense to $(X,x)$ if, for each $R>0$, the sequence of closed balls $\{B_{R}(x_{i})\}$ converges to $B_{R}(x)$ in the Gromov–Hausdorff sense. The following theorem gives an alternative characterization of the tangent cone of an Alexandrov space to the one in Section 2.2.
Theorem 2.8 ([9, Theorem 7.8.1]).
Let $X$ be an Alexandrov space, fix $x\in X$, and let $T_{x}X=C\Sigma_{x}(X)$ be the tangent cone to $X$ at $x$. Then the pointed metric spaces $(\lambda X,x)$ converge in the pointed Gromov-Hausdorff sense to $(T_{x}X,0)$ as $\lambda\to\infty$.
A map $f\colon(X,d_{X})\to(Y,d_{Y})$ between metric spaces is Lipschitz if there exists a real number $K>0$ such that $d_{Y}(f(x_{1}),f(x_{2}))\leq Kd_{X}(x_{1},x_{2})$ for any $x_{1},x_{2}\in X$. Let $X$, and $Y$ be Alexandrov spaces, $U$ an open subset of $X$, and fix $x\in U$. Let $f\colon U\to Y$ be a Lipschitz map and let $\{r_{j}\}$ be a sequence of positive real numbers such that $r_{j}\to 0$ as $j\to\infty$. Then there exists a limit map
$f_{(r_{j})}=\lim_{\omega}(f_{j})\colon T_{x}X\to T_{f(x)}Y,$
where $f_{j}$ is the rescaled function
$f_{j}=f\colon\left(\frac{1}{r_{j}}X,x\right)\to\left(\frac{1}{r_{j}}Y,f(x)\right).$
We say that $f$ is differentiable at $x$, if the map $f_{(r_{j})}$ is independent of the sequence $\{r_{j}\}$. We call this uniquely defined Lipschitz function the differential of $f$ at $x$, and denote it by $df_{x}\colon T_{x}X\to T_{f(x)}Y$. If the differential of $f$ exists at every point $x\in U$, then we say that $f$ is differentiable. If $f$ is differentiable at $x\in U$, then its differential $df_{x}$ is homogeneous. That is, $df_{x}(tv)=tdf(v)$ for each real number $t\geq 0$ and each $v\in T_{x}X$.
The following proposition shows that submetries may transfer geometric properties from the total space to the base space.
Proposition 2.9 ([29, Proposition 4.4]).
Given a submetry $p\colon X\to Y$, the space $Y$ is an Alexandrov space when $X$ is an Alexandrov space.
Moreover, a submetry $p\colon X\to Y$ from an Alexandrov space has a well defined differential.
Proposition 2.10 ([29, Proposition 5.1]).
Let $p\colon X\to Y$ be a submetry from an Alexandrov space $X$ to $Y$ a metric space. Then $p$ is differentiable and each differential $df_{x}\colon T_{x}X\to T_{p(x)}Y$ is a homogeneous submetry.
Let $p\colon X\to Y$ be a submetry from an Alexandrov space $X$ to $Y$ a metric space. A vector $v\in T_{x}X$ is vertical if $df(v)=0$. We call $v$ horizontal if $|Df(v)|=|v|$. From the homogeneity of the differential, one may show that the set of vertical vectors forms a subcone $CV_{x}$ of $T_{x}X$, where $V_{x}\subset\Sigma_{x}X\subset T_{x}X$. Similarly, the set of horizontal vectors forms a subcone $CH_{x}$ of $T_{x}X$, where $H_{x}\subset\Sigma_{x}X\subset T_{x}X$.
A submetry $p\colon X\to Y$ between arbitrary Alexandrov spaces is regular at a point $x\in X$ if the homogeneous submetry $dp_{x}\colon T_{x}X\to T_{p(x)}Y$ is regular. We call the leaf $L_{x}$ a regular leaf. Similarly, $p\colon X\to Y$ is (almost) spanning at $x\in X$ if $dp_{x}\colon T_{x}X\to T_{p(x)}Y$ is (almost) spanning. A point $x\in X$ is round if $\Sigma_{x}(X)$ is a round space. The following corollary follows immediately from Proposition 2.7.
Corollary 2.11.
A submetry $p\colon X\to Y$ from an Alexandrov space $X$ is regular over each round point $y\in Y$.
The following theorem tells us that every Alexandrov space has a dense subset of points whose space of directions is a unit round sphere, and are therefore round points (in the sense defined before Proposition 2.4).
Theorem 2.12 ([8, Theorem 10.8.5] and [8, Corollary 10.9.13]).
If $X$ is an $n$-dimensional Alexandrov space, then the set of points whose space of directions is a unit round sphere (or, equivalently, whose tangent cone is Euclidean space) is dense in $X$.
Recall that, given a submetry $p\colon X\to Y$ between metric spaces, the intrinsic metric on the fiber $L=p^{-1}(y)$ is locally Lipschitz-equivalent to the induced metric on $L$ (see Theorem 2.3). Thus, when $X$ is an Alexandrov space and $p$ is a proper submetry, any fiber $L$ is an Alexandrov space.
The following proposition allows us to identify, for any $x\in X$, the space of vertical directions $V_{x}\subset\Sigma_{x}X$ with $\Sigma_{x}(L)$, the space of directions of the leaf $L$ trough $x$.
Proposition 2.13 ([29, Proposition 5.2]).
Let $p\colon X\to Y$ be a proper submetry between Alexandrov spaces. Then, for any fiber $L=p^{-1}(y)$ and any point $x\in L$, the tangent cone $T_{x}L$ is the cone $CV_{x}$.
Let $I$ be a closed interval. Given $f\colon X\to Y$ and $\alpha\colon I\to Y$, a curve $\tilde{\alpha}\colon I\to X$ is a lift of $\alpha$ if $f\circ\tilde{\alpha}=\alpha$.
Lemma 2.14 (Existence of geodesic lifts: [25, Lemma 1] and [4, Lemma 2.1]).
Let $f\colon X\to Y$ be a submetry between Alexandrov spaces. For any geodesic $\alpha\colon[0,a]\to Y$, and any $x\in f^{-1}(\alpha(0))$, there is a geodesic lift $\tilde{\alpha}\colon[0,a]\to X$ of $\alpha$ starting at $x$, with length equal to that of $\alpha$. Moreover, if $\alpha$ is the geodesic $\alpha$ between $\alpha(0)$ and $\alpha(a)$ is unique, then the lift $\tilde{\alpha}$ is unique.
2.5. Holonomy maps of a submetry between Alexandrov spaces
Let $f\colon X\to Y$ be a submetry between Alexandrov spaces. Assume we have a geodesic $\alpha\colon[0,b]\to Y$ with $\alpha(0)=y$, $\alpha(b)=z$, and initial direction $w\in\Sigma_{y}Y$. By [29, Lemma 5.4], for $x\in f^{-1}(y)$ fixed, for each $h\in H_{x}$ with $df_{x}(h)=w$, there exists a unique geodesic $\tilde{\alpha}_{h}\colon[0,b]\to X$ starting at $x$ with initial direction $h$, such that $f\circ\tilde{\alpha}_{h}=\alpha$. If there exists a unique direction $\tilde{w}\in H_{x}$ such that $df_{x}(\tilde{w})=w$, then we have a well-defined map $p_{\alpha}\colon f^{-1}(y)\to f^{-1}(z)$ given by
$p_{\alpha}(x)=\alpha_{\tilde{w}}(b).$
This map is called the holonomy map of $\alpha$.
Consider a geodesic $\alpha\colon[0,b]\to Y$ such that, for each $x\in f^{-1}(\alpha(0))$, the initial direction of $\alpha$ has a unique lift in $H_{x}$ and $p_{\alpha}$ is well defined. As observed in [29, Section 7.3], for $\tilde{z}\in f^{-1}(\alpha(b))$, the set $p^{-1}_{\alpha}(\tilde{z})$ corresponds to the endpoints of all geodesics of $X$ starting at $\tilde{z}$ which are horizontal lifts of the geodesic $\alpha^{-1}(t)=\alpha(b-t)$ and whose initial direction is a horizontal lift of the initial direction of $\alpha^{-1}$.
Proposition 2.15.
Let $f\colon X\to Y$ be a submetry between Alexandrov spaces and suppose $\alpha\colon[0,b]\to Y$ is a geodesic such that for each $x\in f^{-1}(\alpha(0))$, its initial direction has a unique lift in $H_{x}$ and the holonomy map $p_{\alpha}$ is well defined. Then the holonomy map $p_{\alpha}$ is continuous and surjective.
Proof.
We first prove the surjectivity of the holonomy map $p_{\alpha}$. Consider $z\in f^{-1}(\alpha(b))$. Then, for the geodesic $\alpha^{-1}(t)=\alpha(b-t)$, we have a unique horizontal lift $\widetilde{\alpha^{-1}}$ in $X$. We consider $x=\widetilde{\alpha^{-1}}(b)$, and observe that $f(x)=\alpha(0)$ by construction. The geodesic $(\widetilde{\alpha^{-1}})^{-1}(t)=\widetilde{\alpha^{-1}}(b-t)$ is a horizontal lift of $\alpha$ starting at $x$. Since the horizontal lifts of $\alpha$ are unique by hypothesis, we conclude that $p_{\alpha}(x)=\widetilde{\alpha^{-1}}(b-b)=z$.
We now prove that $p_{\alpha}$ is continuous. Fix $x\in f^{-1}(\alpha(0))$. Consider a sequence $\{x_{i}\}_{i\in\mathbb{N}}\subset f^{-1}(\alpha(0))$ converging to $x$ under the induced metric of $f^{-1}(\alpha(0))$. Since the intrinsic metric is Lipschitz equivalent to the induced metric, we have $d_{X}(x_{i},x)\to 0$ as $i\to\infty$. Denote by $\tilde{\alpha}_{i}$ and $\tilde{\alpha}$ the unique horizontal lifts of $\alpha$ passing through $x_{i}$ and $x$, respectively. By construction, for $t\in[0,b]$, we have $d_{X}(x_{i},\tilde{\alpha}_{i}(t))=d_{Y}(\alpha(0),\alpha(t))=d_{X}(\tilde{\alpha}(t),x)$. Thus, by the continuity of the metric,
$t=d_{X}(\tilde{\alpha}(t),x)$ $=d_{Y}(\alpha(0),\alpha(t))$
$=\lim_{i\to\infty}d_{X}(x_{i},\tilde{\alpha}_{i}(t))$
$=d_{X}\left(\lim_{i\to\infty}x_{i},\lim_{i\to\infty}\tilde{\alpha}_{i}(t)\right)$
$=d_{X}(x,\lim_{i\to\infty}\tilde{\alpha}_{i}(t)).$
This implies that the curve $\beta(t)=\lim_{i\to\infty}\tilde{\alpha}_{i}(t)$ is well defined. By the continuity of $f$,
$(f\circ\beta)(t)=f\left(\lim_{i\to\infty}\tilde{\alpha}_{i}(t)\right)=\lim_{i\to\infty}f(\tilde{\alpha}_{i}(t))=\lim_{i\to\infty}\alpha(t)=\alpha(t).$
Moreover, for any $s,t\in[0,b]$, we have
$d_{X}(\tilde{\alpha}(t),\tilde{\alpha}(s))$ $=|s-t|$
$=d_{Y}(\alpha(s),\alpha(t))$
$=\lim_{i\to\infty}d_{X}(\tilde{\alpha}_{i}(s),\tilde{\alpha}_{i}(t))$
$=d_{X}\left(\lim_{i\to\infty}\tilde{\alpha}_{i}(s),\lim_{i\to\infty}\tilde{\alpha}_{i}(t)\right)$
$=d_{X}(\beta(s),\beta(t).$
From this we conclude that $\beta$ is a geodesic in $X$ and it is a horizontal lift of $\alpha$ starting at $\beta(0)=\lim_{i\to\infty}\tilde{\alpha}_{i}(0)=\lim_{i\to\infty}x_{i}=x$. By the uniqueness of the horizontal lift of $\alpha$, we conclude that $\tilde{\alpha}=\beta$. In particular,
$\lim_{i\to\infty}p_{\alpha}(x_{i})=\lim_{i\to\infty}\tilde{\alpha}_{i}(b)=\beta(b)=\tilde{\alpha}(b)=p_{\alpha}(x).$
Thus, $p_{\alpha}$ is continuous. ∎
Corollary 2.16.
Let $f\colon X\to Y$ be a submetry between Alexandrov spaces. If $\alpha\colon[0,b]\to Y$ is a geodesic whose endpoints are regular points, then the holonomy map $p_{\alpha}$ is well defined and is a homeomorphism.
Proof.
Let $y=\alpha(0)$ and $z=\alpha(b)$. By definition, for $x\in f^{-1}(y)$ we have $df_{x}\colon H_{x}\to\Sigma_{y}Y$ is an isometry. Thus, the holonomy map $p_{\alpha}$ is well defined. Moreover, for $\tilde{z}\in f^{-1}(z)$, the holonomy map $p_{\alpha^{-1}}$ is also well defined for the geodesic $\alpha^{-1}(t)=\alpha(b-t)$. Additionally, $p_{\alpha^{-1}}(p_{\alpha}(x))=x$ by construction. By Proposition 2.15, both $p_{\alpha}$ and $p_{\alpha^{-1}}$ are continuous. Thus, the conclusion follows. ∎
The following assertion is now an immediate consequence of Corollary 2.16.
Corollary 2.17.
Let $f\colon X\to Y$ be a submetry between Alexandrov spaces. Then the regular fibers of $f$ are homeomorphic.
2.6. Isometry groups of Alexandrov spaces
As for Riemannian manifolds, the isometry group of an Alexandrov space admits a Lie group structure. Additionally, there is an optimal upper bound on the dimension of the isometry group, and a rigidity result holds when this bound is attained.
Theorem 2.18 ([17, Theorem 1.1] and [18, Theorems 3.1 and 4.1]).
Let $(X,d)$ be an $n$-dimensional Alexandrov space. Then the following assertions hold:
  1. (1)
    The group $G$ of isometries of $X$ is a Lie group and is compact if $X$ is compact.
  2. (2)
    The dimension of $G$ satisfies $\dim(G)\leqslant n(n+1)/2$.
  3. (3)
    If $\dim(G)=n(n+1)/2$, then $X$ is a Riemannian manifold isometric to one of the following spaces of constant sectional curvature: the $n$-dimensional Euclidean space, an $n$-dimensional round sphere, an $n$-dimensional round real projective space, or an $n$-dimensional hyperbolic space with constant negative sectional curvature.
Remark 2.19.
The conclusions of Theorem 2.18 (1) also hold for a complete length space $(X,d)$ which is locally an Alexandrov space. That is, for each $x\in X$, there exist a neighborhood $U_{x}\subset X$ and a real number $k_{x}$ such that $(U_{x},{\left.\kern-1.2ptd\right|_{U_{x}}})$ is an Alexandrov space with $\mathrm{curv}\geq k$.
Let $g\colon X\to X$ be an isometry of an Alexandrov space $X$, and suppose that there exists $x\in X$ such that $g(x)=x$. Then the differential $dg_{x}\colon T_{x}X\to T_{x}X$ induces an isometry on the space of directions $\Sigma_{x}(X)\subset T_{x}X=C\Sigma_{x}(X)$. Let $G\subset\operatorname{Isom}(X)$ be a group acting on $X$ isometrically. Given $x\in X$, we define the orbit of $G$ through $x$ as $G(x)=\{g\cdot x\mid g\in G\}$. The isotropy subgroup at $x$ is the group $G_{x}=\{g\in G\mid g\cdot x=x\}$. The isotropy representation of $G_{x}$ into $\operatorname{Isom}(\Sigma_{x}(X))$ is given by setting $g\cdot v=dg_{x}(v)$ for $g\in G_{x}$ and $v\in\Sigma_{x}(X)$. The action of $G$ on $X$ is effective if $\cap_{x\in X}G_{x}=\{e\}$. The following lemma implies that, if $G$ acts effectively on $X$, then, for any $x\in X$, the action of $G_{x}$ on $\Sigma_{x}(X)$ given by the isotropy representation is also effective.
Lemma 2.20 ([18, Lemma 3.2]).
Let $X$ be an Alexandrov space and let $g\colon X\to X$ an isometry. If there is some $x_{0}\in X$ such that $g(x_{0})=x_{0}$ and $dg_{x_{0}}\colon\Sigma_{x_{0}}(X)\to\Sigma_{x_{0}}(X)$ is the identity, then $g(x)=x$ for all $x\in X$.
2.7. Foliated homeomorphisms
To conclude this section, let us recall some results on foliated homeomorphisms (cf. [32, Section 3]). Let $p\colon X\to Y$ be a quotient map between two topological spaces $X$ and $Y$. Denote by =$\mathcal{F}=\{p^{-1}(y)\mid y\in Y\}$ the partition of $X$ induced by the preimages of $p$. A continuous map $h\colon X\to X$ is an $\mathcal{F}$-map if, for some $L\in\mathcal{F}$, we have $h(L)\subset L^{\prime}$ for some $L^{\prime}\in\mathcal{F}$. Thus, every $\mathcal{F}$-map $h\colon X\to X$ induces a well defined map $\psi(h)\colon Y\to Y$ making the following diagram commute:
${X}$${X}$${Y}$${Y}$$\scriptstyle{p}$$\scriptstyle{h}$$\scriptstyle{p}$$\scriptstyle{\psi(h)}$
We denote by $\operatorname{\mathrm{Homeo}}(X,\mathcal{F})$ the group of all $\mathcal{F}$-homeomorphisms (i.e., all $\mathcal{F}$-maps with an inverse map which is an $\mathcal{F}$-map). We let $\operatorname{\mathrm{Homeo}}(Y)$ be the group of all self-homeomorphisms of $Y$. We have a group homomorphism $\psi\colon\operatorname{\mathrm{Homeo}}(X,\mathcal{F})\to\operatorname{\mathrm{Homeo}}(Y)$. The map $p\colon X\to Y$ admits local cross-sections if, for any $y\in Y$, there exists an open neighborhood $V\subset Y$ and a continuous map $\sigma\colon V\to X$ such that $p\circ\sigma=\mathrm{Id}_{V}$. When $Y$ is locally compact and Hausdorff, the following theorem guarantees the continuity of $\psi$ when we equip both $\operatorname{\mathrm{Homeo}}(X,\mathcal{F})$ and $\operatorname{\mathrm{Homeo}}(Y)$ with the compact-open topology.
Theorem 2.21 ([32, Corollary 3.4]).
Let $X,Y$ be topological spaces with $Y$ locally compact and Hausdorff. If $p\colon X\to Y$ is a quotient map, then each of the following conditions implies that the homomorphism $\psi\colon\operatorname{\mathrm{Homeo}}(X,\mathcal{F})\to\operatorname{\mathrm{Homeo}}(Y)$ is continuous:
  1. (1)
    The map $p$ is proper.
  2. (2)
    The map $p$ is open map and admits local cross-sections.
  3. (3)
    The map $p$ is a locally trivial fibration.
3. Proof of Theorem A
We will prove each item in Theorem A separately. To begin, we prove some auxiliary results on proper metric submetries.
Proposition 3.1.
Let $p\colon X\to Y$ be a proper submetry from a length space $X$ onto a metric space $Y$ and set =$\mathcal{F}=\{p^{-1}(y)\mid y\in Y\}$. If the group $\operatorname{Isom}(Y)$ is closed, then the group $\operatorname{Isom}(X,\mathcal{F})$ of foliated isometries is a closed subset of $\operatorname{Isom}(X)$.
Proof.
Let $\{f_{n}\}\subset\operatorname{Isom}(X,\mathcal{F})\subset\operatorname{\mathrm{Homeo}}(X)$ be a sequence of foliated isometries of $X$ and suppose that there exists an isometry $f\in\operatorname{Isom}(X)$ such that $f_{n}\to f$ in the compact-open topology of $\operatorname{\mathrm{Homeo}}(X)$. We will show that $f$ is a foliated map.
Each foliated isometry $f_{n}$ of $X$ induces an isometry $f^{\ast}_{n}$ of $Y$, defined as $f^{\ast}_{n}(x^{\ast})=f^{\ast}_{n}(p(x))=p(f_{n}(x))$, where $x\in X$ is any point in the preimage of $x^{\ast}$ under the submetry $p\colon X\to Y$. Since $f_{n}$ is a foliated map, $f^{\ast}_{n}$ is well defined, i.e., its definition is independent of the choice of $x\in p^{-1}(x^{\ast})$.
Consider the map $\psi\colon\operatorname{\mathrm{Homeo}}(X,\mathcal{F})\to\operatorname{\mathrm{Homeo}}(Y)$ mapping a foliated homeomorphism $h\in\operatorname{\mathrm{Homeo}}(X,\mathcal{F})$ to $\psi(h)=p\circ h$. By Theorem 2.21, $\psi$ is continuous with respect to the compact-open topology. Restricting $\psi$ to the foliated isometries yields a map $\Psi\colon\operatorname{Isom}(X,\mathcal{F})\to\operatorname{Isom}(Y)$ that is also continuous with respect to the compact-open topology.
Fix $x,x^{\prime}\in X$ such that $p(x)=p(x^{\prime})$. By construction, $f_{n}(x)\to f(x)$ and $f_{n}(x^{\prime})\to f(x^{\prime})$ as $n\to\infty$. Since $p$ is continuous we have that $p(f_{n}(x))\to p(f(x))$ and $p(f_{n}(x^{\prime}))\to p(f(x^{\prime}))$ as $n\to\infty$. Given that $f_{n}\in\operatorname{\mathrm{Homeo}}(X,\mathcal{F})$, then we have that $p(f_{n}(x))=p(f_{n}(x^{\prime}))$. By uniqueness of the limit, we have that $p(f(x))=p(f(x^{\prime}))$. This implies that for $x,x^{\prime}\in X$ such that $p(x)=p(x^{\prime})$ we have $p(f(x))=p(f(x^{\prime}))$, and thus $f\in\operatorname{\mathrm{Homeo}}(X,\mathcal{F})$. Thus, $f$ is a foliated isometry. ∎
Since a closed subgroup of a Lie group is itself a Lie group, Proposition 3.1 yields the following corollary.
Corollary 3.2.
Let $p\colon X\to Y$ be a proper submetry from a length space $X$ to a metric space $Y$ and set =$\mathcal{F}=\{p^{-1}(y)\mid y\in Y\}$. If $\operatorname{Isom}(X)$ admits a Lie group structure with respect to the open-compact topology and $\operatorname{Isom}(Y)$ is closed, then $\operatorname{Isom}(X,\mathcal{F})$ is a Lie group.
3.1. Proof of Theorem A (i)
Let $(M,\mathcal{F})$ be a singular Riemannian foliation with closed leaves on a complete Riemannian manifold and equip the leaf space $M/\mathcal{F}$ with the leaf-projection metric. As recalled in Section 2.3, the leaf-projection map $\pi\colon M\to M/\mathcal{F}$ is a submetry.
Lemma 3.3.
The leaf-projection map $\pi\colon M\to M/\mathcal{F}$ is proper.
Proof.
Let $K$ be a compact subset of the leaf space $M/\mathcal{F}$. Since $M/\mathcal{F}$ is a complete metric space, $K$ is totally bounded. Thus, given $\varepsilon>0$, we may cover $K$ with a finite number of closed balls of radius $\varepsilon$. By the slice theorem for singular Riemannian foliations (see [34, Theorem A]), we can choose $\varepsilon$ in such a way that each $\varepsilon$-ball in the cover corresponds to a closed tubular neighborhood of some leaf. Thus, the preimage of $K$ is contained in a finite union of compact sets, and is therefore compact. Hence, the leaf-projection map is proper. ∎
Lemma 3.4.
The group of isometries of $M/\mathcal{F}$ is a closed subgroup of $\operatorname{\mathrm{Homeo}}(M/\mathcal{F})$.
Proof.
Let $\{f_{n}^{\ast}\}\subset\operatorname{Isom}(M/\mathcal{F})$ be a sequence with limit $f^{\ast}\in\operatorname{\mathrm{Homeo}}(M/\mathcal{F})$. Let us show that $f^{\ast}\in\operatorname{Isom}(M/\mathcal{F})$. Consider two distinct points $p^{\ast}$ and $q^{\ast}$ in $M/\mathcal{F}$. Applying twice the triangle inequality, we obtain
$d_{M/\mathcal{F}}(f^{\ast}(p^{\ast}),f^{\ast}(q^{\ast}))\leq d_{M/\mathcal{F}}(f^{\ast}(p^{\ast}),f_{n}^{\ast}(p^{\ast}))+d_{M/\mathcal{F}}(f_{n}^{\ast}(p^{\ast}),f_{n}^{\ast}(q^{\ast}))+d_{M/\mathcal{F}}(f_{n}^{\ast}(q^{\ast}),f^{\ast}(q^{\ast})).$
Taking $\varepsilon>0$, assuming that $n$ is sufficiently large, and using the fact that each $f^{\ast}_{n}$ is an isometry of $M/\mathcal{F}$, we obtain that
$d_{M/\mathcal{F}}(f^{\ast}(p^{\ast}),f^{\ast}(q^{\ast}))$ $\leq d_{M/\mathcal{F}}(f_{n}^{\ast}(p^{\ast}),f_{n}^{\ast}(q^{\ast}))+2\varepsilon$
$=d_{M/\mathcal{F}}(p^{\ast},q^{\ast})+2\varepsilon.$
Thus,
(3.1) $d_{M/\mathcal{F}}(f^{\ast}(p^{\ast}),f^{\ast}(q^{\ast}))-d_{M/\mathcal{F}}(p^{\ast},q^{\ast})\leq 2\varepsilon.$
Similarly, using again the fact that $f_{n}^{*}$ is an isometry of $M/\mathcal{F}$ and applying twice the triangle inequality, we obtain
$d_{M/\mathcal{F}}(p^{\ast},q^{\ast})\leq d_{M/\mathcal{F}}(f_{n}^{\ast}(p^{\ast}),f^{\ast}(p^{\ast}))+d_{M/\mathcal{F}}(f^{\ast}(p^{\ast}),f^{\ast}(q^{\ast}))+d_{M/\mathcal{F}}(f^{\ast}(q^{\ast}),f_{n}^{\ast}(q^{\ast})).$
Assuming again that $n$ is sufficiently large, we obtain that
$d_{M/\mathcal{F}}(p^{\ast},q^{\ast})\leq d_{M/\mathcal{F}}(f^{\ast}(p^{\ast}),f^{\ast}(q^{\ast}))+2\varepsilon.$
Thus,
(3.2) $d_{M/\mathcal{F}}(p^{\ast},q^{\ast})-d_{M/\mathcal{F}}(f^{\ast}(p^{\ast}),f^{\ast}(q^{\ast}))\leq 2\varepsilon.$
Inequalities (3.1) and (3.2) imply that, for any $\varepsilon>0$,
$|d_{M/\mathcal{F}}(f^{\ast}(p^{\ast}),f^{\ast}(q^{\ast}))-d_{M/\mathcal{F}}(p^{\ast},q^{\ast})|\leq 2\varepsilon.$
This implies that $d_{M/\mathcal{F}}(f^{\ast}(p^{\ast}),f^{\ast}(q^{\ast}))=d_{M/\mathcal{F}}(p^{\ast},q^{\ast})$. Since $p^{\ast}$ and $q^{\ast}$ are arbitrary, we conclude that $f^{\ast}$ is an isometry. ∎
With the preceding observations in hand, we may now prove the first item in Theorem A.
Proof of Theorem A (i).
First, recall that the Riemannian metric of $M$ induces a metric $d^{\ast}$ on $M/\mathcal{F}$, making it a complete length space $(M/\mathcal{F},d^{\ast})$. Moreover, $(M/\mathcal{F},d^{\ast})$ is locally an Alexandrov space, as pointed out in Remark 2.19 (see [30, pp. 77–78]). Thus, by Lemma 3.3, the map $\pi\colon M\to M/\mathcal{F}$ is a proper submetry, and by Lemma 3.4, the group $\operatorname{Isom}(M/\mathcal{F})$ is closed. Moreover, by Theorem 2.18 (1), since $(M/\mathcal{F})$ is locally an Alexandrov space, $\operatorname{Isom}(M/\mathcal{F})$ is a Lie group. The desired conclusion now follows from Corollary 3.2. ∎
3.2. Proof of Theorem A (ii)
We first recall the following useful fact, which we will use in the proof of the theorem.
Lemma 3.5.
Let $f,g\colon X\to Y$ be continuous functions between topological spaces. If $Y$ is Hausdorff and there exists a dense subset $D\subset X$ of $X$ such that $f|_{D}=g|_{D}$, then $f=g$.
Proof.
We prove the statement by contradiction. Assume that there exists $x_{0}\in X$ such that $f(x_{0})\neq g(x_{0})$. Since $Y$ is Hausdorff, there exist open neighborhoods $U$ and $V$ of $f(x_{0})$ and $g(x_{0})$, respectively, with $U\cap V=\varnothing$. Observe now that $W=f^{-1}(U)\cap g^{-1}(V)\subset X$ is a non-empty open subset. Since $D$ is dense in $X$, there exists $y\in D\cap W$ and, by hypothesis, $f(y)=g(y)$. Since $y\in W$, $f(y)\in U$ and $g(y)\in V$. This implies that $V\cap U\neq\varnothing$, which is a contradiction. ∎
Proof of Theorem A (ii).
Recall that $M_{\mathrm{prin}}\subset M$, the set of principal leaves, is an open, connected, and dense subset of $M$ (see Section 2.1). The quotient $M_{\mathrm{prin}}/\mathcal{F}$ is a Riemannian manifold. Hence, by the Myers–Steenrod theorem, $\operatorname{Isom}(M_{\mathrm{prin}}/\mathcal{F})$ is a Lie group. Moreover, $\operatorname{Isom}(M_{\mathrm{prin}}/\mathcal{F})$ is closed (for example, by the proof of Lemma 3.4). In addition, the projection map $\pi\colon M_{\mathrm{prin}}\to M_{\mathrm{prin}}/\mathcal{F}$ is proper. These observations, along with Proposition 3.1, imply that $\operatorname{Isom}(M_{\mathrm{prin}},\mathcal{F})$ is a closed subgroup of $\operatorname{Isom}(M_{\mathrm{prin}})$. Now, since $M_{\mathrm{prin}}$ is a smooth submanifold of $M$, $\operatorname{Isom}(M_{\mathrm{prin}})$ is a Lie group. Therefore, $\operatorname{Isom}(M_{\mathrm{prin}},\mathcal{F})$ is a Lie group.
Denote the identity map on $M$ by $\mathrm{Id}_{M}$ and its restriction to $M_{\mathrm{prin}}$ by $\mathrm{Id}_{M_{\mathrm{prin}}}$. Consider the homomorphism $\Phi\colon\operatorname{Isom}(M,\mathcal{F})\to\operatorname{Isom}(M_{\mathrm{prin}},\mathcal{F})$ given by $\Phi(f)=f|_{M_{\mathrm{prin}}}$. Note that $\Phi$ is injective. Assume that $\Phi(f)=\mathrm{Id}_{M_{\mathrm{prin}}}$. Since $M_{\mathrm{prin}}$ is dense and $M$ is Hausdorff, $f=\mathrm{Id}_{M}$ by Lemma 3.5. Furthermore, the map $\Phi$ is continuous and, therefore, smooth by [26, Corollary 3.50]. Hence, $\Phi$ is a Lie group homomorphism.
Fix $v\in T_{\mathrm{Id}_{M}}\operatorname{Isom}(M,\mathcal{F})$ and assume that $\Phi_{\ast_{\mathrm{Id}_{M}}}(v)=0$. We will show that $v=0$, which implies that $\Phi_{\ast_{\mathrm{Id}_{M}}}$ is injective. Since $\Phi$ is a Lie group homomorphism and $\Phi(\mathrm{Id}_{M})=\mathrm{Id}_{M_{\mathrm{prin}}}$, we have
(3.3) $\Phi(\exp_{\mathrm{Id}_{M}}(v))=\exp_{\mathrm{Id}_{M_{\mathrm{prin}}}}(\Phi_{\ast_{\mathrm{Id}_{M}}}(v))=\mathrm{Id}_{M_{\mathrm{prin}}}.$
By definition, $\Phi(\exp_{\mathrm{Id}_{M}}(v))=\exp_{\mathrm{Id}_{M}}(v)|_{M_{\mathrm{prin}}}$. Hence, equation (3.3) implies that the isometry $\exp_{\mathrm{Id}_{M}}(v)\colon M\to M$ restricted to the open and dense subset $M_{\mathrm{prin}}\subset M$ is the identity map, i.e., $\exp_{\mathrm{Id}_{M}}(v)|_{M_{\mathrm{prin}}}=\mathrm{Id}_{M_{\mathrm{prin}}}=\mathrm{Id}_{M}|_{M_{\mathrm{prin}}}$. Then, by Lemma 3.5, $\exp_{\mathrm{Id}_{M}}(v)=\mathrm{Id}_{M}$. This implies that $v=0$. Therefore, $\Phi_{\ast_{\mathrm{Id}_{M}}}$ is injective and it follows that
$\dim(\operatorname{Isom}(M,\mathcal{F}))\leq\dim(\operatorname{Isom}(M_{\mathrm{prin}},\mathcal{F})).$
Thus, to prove inequality (a), it suffices to show that
(3.4) $\dim(\operatorname{Isom}(M_{\mathrm{prin}},\mathcal{F}))\leq\frac{k(k+1)}{2}+\frac{(n-k)(n-k+1)}{2}.$
The projection $\pi\colon M_{\mathrm{prin}}\to M_{\mathrm{prin}}/\mathcal{F}$ is a fiber bundle with $(n-k)$-dimensional fibers (see [34, Theorem A] and [10, p. 41]). Hence, $M_{\mathrm{prin}}/\mathcal{F}\subset M/\mathcal{F}$ is a $k$-dimensional smooth manifold and inherits a Riemannian metric from $M_{\mathrm{prin}}$, making the map $\pi\colon M_{\mathrm{prin}}\to M_{\mathrm{prin}}/\mathcal{F}$ a Riemannian submersion. By the classical Myers–Steenrod theorem [37], both $\operatorname{Isom}(M_{\mathrm{prin}})$ and $\operatorname{Isom}(M_{\mathrm{prin}}/\mathcal{F})$ are Lie groups. Moreover, by [28, II Theorem 3.1],
(3.5) $\dim(\operatorname{Isom}(M_{\mathrm{prin}}/\mathcal{F}))\leq\frac{k(k+1)}{2}.$
Consider now the group homomorphism $\Psi\colon\operatorname{Isom}(M_{\mathrm{prin}},\mathcal{F})\to\operatorname{\mathrm{Homeo}}(M_{\mathrm{prin}}/\mathcal{F})$ given by
$\Psi(f)(p^{\ast})=(\pi\circ f)(q),$
where $q\in\pi^{-1}(p^{\ast})$. Let $G=\ker(\Psi)$. Since the leaves of $\mathcal{F}$ are closed, by [10] and [34], for any principal leaf $L\in\mathcal{F}$ which has dimension $(n-k)$, a sufficiently small tubular neighborhood $\mathrm{Tub}(L)$ of $L$ is foliated diffeomorphic to
$\mathrm{Tub}(L)\cong\left(L\times D^{k},\{L\times\{x\}\mid x\in D^{k}\}\right),$
where $D^{k}$ denotes a $k$-dimensional disk. We point out that $\mathrm{Tub}(L)^{\ast}\cong D^{k}$. Thus, by definition, the group $G$ induces a trivial action on $D^{k}$, and $G$ acts effectively on the tubular neighborhood; otherwise, the group $\operatorname{Isom}(M_{\mathrm{prin}},\mathcal{F})$ would not act effectively. In this way, we get an effective action of $G$ on $L$. Moreover, this action is by isometries with respect to the induced metric of $M$ on $L$. Thus, by [28, II Theorem 3.1],
(3.6) $\dim(G)\leqslant\frac{(n-k)(n-k+1)}{2}.$
Since $\Psi\colon\operatorname{Isom}(M_{\mathrm{prin}},\mathcal{F})\to\operatorname{Isom}(M_{\mathrm{prin}}/\mathcal{F})$ is a continuous Lie group homomorphism, $\Psi$ is smooth. Moreover, $\Psi(\operatorname{Isom}(M_{\mathrm{prin}},\mathcal{F}))$ is a Lie subgroup of $\operatorname{Isom}(M_{\mathrm{prin}}/\mathcal{F})$. Then, by the rank-nullity theorem and inequalities (3.5) and (3.6),
$\dim(\operatorname{Isom}(M_{\mathrm{prin}},\mathcal{F}))$ $=\dim(G)+\dim(\Psi(\operatorname{Isom}(M_{\mathrm{prin}},\mathcal{F})))$
$\leqslant\dim(G)+\dim(\operatorname{Isom}(M_{\mathrm{prin}}/\mathcal{F}))$
$\leqslant\frac{(n-k)(n-k+1)}{2}+\frac{k(k+1)}{2}.$
This last series of inequalities verifies inequality (3.4), which in turn implies inequality (a). ∎
Remark 3.6.
Observe that $f\in\ker(\Psi\colon\operatorname{Isom}(M,\mathcal{F})\to\operatorname{Isom}(M^{\ast}))$ if and only if $f(L)\subset L$ for any leaf $L\in\mathcal{F}$. Now, for the Lie group homomorphism $\Phi\colon\operatorname{Isom}(M,\mathcal{F})\to\operatorname{Isom}(M_{\mathrm{prin}},\mathcal{F})$ in the proof of Theorem A (ii), if $f\in\ker\left(\Psi\colon\operatorname{Isom}(M,\mathcal{F})\to\operatorname{Isom}(M^{\ast})\right)$, then $\Phi(f)(L)=L$ for any leaf $L\subset M_{\mathrm{prin}}$. This implies that $\Phi(f)\in\ker(\Psi\colon\operatorname{Isom}(M_{\mathrm{prin}},\mathcal{F})\to\operatorname{Isom}(M_{\mathrm{prin}}^{\ast}))$. Thus, we conclude that
$\Phi(\ker(\Psi\colon\operatorname{Isom}(M,\mathcal{F})\to\operatorname{Isom}(M^{\ast}))\subset\ker(\Psi\colon\operatorname{Isom}(M_{\mathrm{prin}},\mathcal{F})\to\operatorname{Isom}(M_{\mathrm{prin}}^{\ast})).$
This, together with the fact that $\Phi_{\ast_{\mathrm{Id}_{M}}}$ is injective, implies that
$\dim(\ker(\Psi\colon\operatorname{Isom}(M,\mathcal{F})\to\operatorname{Isom}(M^{\ast})))\leq\frac{k(k+1)}{2}.$
With this observation in hand, we may now finish the proof of Theorem A.
3.3. Proof of Theorem A (iii)
Consider the Lie group homomorphism $\Psi\colon\operatorname{Isom}(M,\mathcal{F})\to\operatorname{Isom}(M/\mathcal{F})$. Since $\operatorname{Isom}(M,\mathcal{F})$ has maximal dimension, Remark 3.6 implies that the kernel $G=\ker\Psi$ has dimension equal to $k(k+1)/2$. Moreover, since $G$ acts effectively by isometries on any leaf, we conclude, by [28, II Theorem 3.1], that all the leaves are isometric to a round sphere $S^{k}$ or a round real projective space $\mathbb{R}P^{k}$. This implies that the foliation $\mathcal{F}$ is regular.
Equality in the dimension of $\operatorname{Isom}(M,\mathcal{F})$ implies that the dimension of $\operatorname{Isom}(M/\mathcal{F})$ is equal to $(n-k)(n-k+1)/2$. The proof of Theorem A (ii) in turn implies that $\dim(\operatorname{Isom}(M_{\mathrm{prin}}/\mathcal{F})=(n-k)(n-k+1)/2$. Since $M_{\mathrm{prin}}/\mathcal{F}$ is a Riemannian manifold, we conclude, by [28, II Theorem 3.1], that $M_{\mathrm{prin}}/\mathcal{F}$ is isometric to one of the following: an $(n-k)$-dimensional Euclidean space, an $(n-k)$-dimensional hyperbolic space with constant negative curvature, an $(n-k)$-dimensional round real projective space, or an $(n-k)$-dimensional round sphere. Therefore, $M_{\mathrm{prin}}/\mathcal{F}$ is complete. Since $M^{\ast}$ is the metric completion of $M_{\mathrm{prin}}/\mathcal{F}$, it follows that $M^{\ast}=M_{\mathrm{prin}}/\mathcal{F}$. Recall that over the principal leaves of the singular Riemannian foliation $\mathcal{F}$, the map $\pi\colon M_{\mathrm{prin}}\to M_{\mathrm{prin}}/\mathcal{F}$ is a smooth Riemannian submersion. This implies that $\pi\colon M\to M/\mathcal{F}$ is a fiber bundle, and thus the desired conclusion follows. ∎
4. Proof of Theorem B
We will prove each item in Theorem B separately. Throughout this section, we let $\pi\colon X\to Y$ be a proper submetry between Alexandrov spaces and set =$\mathcal{F}=\{\pi^{-1}(y)\mid y\in Y\}$.
4.1. Proof of Theorem B (i)
Since $Y$ is an Alexandrov space, Theorem 2.18 (1) implies that $\operatorname{Isom}(Y)$ is a closed Lie group. Then, by Corollary 3.2, $\operatorname{Isom}(X,\mathcal{F})$ is a Lie group. Moreover, when $X$ is compact, $Y$ is also compact and, hence, $\operatorname{Isom}(X)$ and $\operatorname{Isom}(Y)$ are compact, and thus closed. Consider the map $\Psi\colon\operatorname{Isom}(X,\mathcal{F})\to\operatorname{Isom}(Y)$ given by $\Psi(f)(y)=(\pi\circ f)(x)$ for any $x\in\pi^{-1}(y)$ (cf. Section 2.7). By Proposition 3.1, $\operatorname{Isom}(X,\mathcal{F})$ is closed in $\operatorname{Isom}(X)$, since $\operatorname{Isom}(Y)$ is closed. Thus, we conclude that $\operatorname{Isom}(X,\mathcal{F})$ is compact. ∎
4.2. Proof of Theorem B (ii)
We now compute the upper bound on the dimension of the Lie group $\operatorname{Isom}(X,\mathcal{F})$. Suppose that $X$ has dimension $n\geq 1$ and $Y$ has dimension $0\leqslant k\leqslant n$, so that the regular leaves of the foliation $\mathcal{F}$ have dimension $n-k$. We will show that
(4.1) $\dim(\operatorname{Isom}(X,\mathcal{F}))\leqslant\frac{k(k+1)}{2}+\frac{(n-k)(n-k+1)}{2}.$
Recall that a point $y\in Y$ is regular with respect to $\pi$ if $\pi$ is regular at each $x\in\pi^{-1}(y)$ (see [29, Definition 6.5]). The set of regular points $Y_{\mathrm{reg}}=\{y\in Y\mid y\mbox{ is regular with respec to }\pi\}$ contains the open and dense subset of thick points (see [29, End of Section 6]). Consequently, the set $X_{\mathrm{reg}}=\{x\in X\mid x\mbox{ is regular}\}$ contains an open and dense subset. Let $R_{X}=\{x\in X\mid\Sigma_{x}X\mbox{ is isometric to the unit round }S^{n-1}\}$. By Theorem 2.12, $R_{X}$ is dense. Hence, $R_{X}\cap X_{\mathrm{reg}}\neq\varnothing$.
Fix $\bar{x}\in R_{X}\cap X_{\mathrm{reg}}$. We may choose $\bar{x}$ so that $\Sigma_{\pi(\bar{x})}Y$ is isometric to the unit round $S^{k}$, since the set of points in $Y$ whose space of directions is isometric to the unit round $S^{k}$ is dense in $Y$ and $Y_{\mathrm{reg}}$ contains an open subset. Since $\bar{x}$ is regular, $d_{\bar{x}}\pi|_{H_{\bar{x}}}\colon H_{\bar{x}}\to\Sigma_{\pi(\bar{x})}Y$ is an isometry (see Section 2.4). Moreover, since $\Sigma_{\pi(\bar{x})}Y$ is round, there exists an isometry between $\Sigma_{\bar{x}}X$ and the spherical join $H_{\bar{x}}\star V_{\bar{x}}$. Since $\Sigma_{\bar{x}}X$ is isometric to the unit round $S^{n-1}$ and $H_{\bar{x}}$ is isometric to the unit round $S^{k-1}$, we conclude that $V_{\bar{x}}$ is isometric to the unit round $S^{n-k-1}$.
Consider now the Lie group homomorphism $\Psi\colon\operatorname{Isom}(X,\mathcal{F})\to\operatorname{Isom}(Y)$ given by $\Psi(f)=\pi\circ f$ (cf. Section 2.7 and Section 3) and set $G=\ker(\Psi)$.
Given $f\in G$, we have $f(\bar{x})\in\pi^{-1}(\pi(\bar{x}))$, allowing us to identify the differential $d_{\bar{x}}f\colon\Sigma_{\bar{x}}X\to\Sigma_{f(\bar{x})}X$ (wich is an isometry) with an isometry
$(d_{\bar{x}}f)_{1}\star(d_{\bar{x}}f)_{2}\colon H_{\bar{x}}\star V_{\bar{x}}\to H_{f(\bar{x})}\star V_{f(\bar{x})}.$
Since $G$ acts on $\pi^{-1}(\pi(\bar{x}))$, we may consider the isotropy group $G_{\bar{x}}$. For $f\in G_{\bar{x}}$, we have then that $(d_{\bar{x}}f)_{1}=\mathrm{Id}_{H_{\bar{x}}}$ and $(d_{\bar{x}}f)_{2}\colon V_{\bar{x}}\to V_{\bar{x}}$ is an isometry. Thus, $G_{\bar{x}}$ acts by isometries on the unit round $S^{n-k-1}$. Moreover, this action is effective. Indeed, if $(d_{\bar{x}}f)_{2}=\mathrm{Id}_{V_{\bar{x}}}$, then $d_{\bar{x}}f=\mathrm{Id}_{\Sigma_{\bar{x}}X}$. Since $f(\bar{x})=\bar{x}$ and $f$ is an isometry of the Alexandrov space $X$, then [18, Lemma 3.2] implies that $f=\mathrm{Id}_{X}$. Hence, by the Myers–Steenrod theorem,
$\dim(G_{\bar{x}})\leq\frac{(n-k-1)(n-k)}{2}.$
Now we consider the orbit $G(\bar{x})\subset\pi^{-1}(\pi(\bar{x}))$. By [29, Korollar 7.5], the Hausdorff dimension of $\pi^{-1}(\pi(\bar{x}))$ is $n-k$. Thus, we conclude that the dimension of $G(\bar{x})=G/G_{\bar{x}}$ is bounded above by $n-k$. It follows then that
(4.2) $\begin{split}\dim(G)=\dim(G(\bar{x}))+\dim(G_{\bar{x}})&\leq n-k+\frac{(n-k-1)(n-k)}{2}\\[5.69046pt] &=\frac{(n-k+1)(n-k)}{2}.\end{split}$
Next, observe that $\Psi(\operatorname{Isom}(X,\mathcal{F}))$ is a Lie subgroup of $\operatorname{Isom}(Y)$. Since $Y$ is an Alexandrov space of dimension $k$, Theorem 2.18 implies that
(4.3) $\dim(\Psi(\operatorname{Isom}(X,\mathcal{F})))\leqslant k(k+1)/2.$
Then, by the first group isomorphism theorem, inequalities (4.2) and (4.3), and recalling that $G=\ker\Psi$,
(4.4) $\begin{split}\dim\left(\operatorname{Isom}(X,\mathcal{F})\right)&=\dim\left(G\right)+\dim\left(\Psi(\operatorname{Isom}(X,\mathcal{F})\right))\\[5.69046pt] &\leq\frac{(n-k)(n-k+1)}{2}+\frac{k(k+1)}{2}.\end{split}$
4.3. Proof of Theorem B (iii)
We now examine the case where equality in inequality (4.4) holds. We first determine the topology of the base space $Y$ of the submetry $\pi\colon X\to Y$.
Proposition 4.1.
Let $\pi\colon X\to Y$ be a proper submetry from an Alexandrov space $X$ of dimension $n$ onto an Alexandrov space $Y$ of dimension $0\leqslant k\leqslant n$. If
$\dim(\operatorname{Isom}(X,\mathcal{F}))=\frac{k(k+1)}{2}+\frac{(n-k)(n-k+1)}{2},$
then $Y$ is isometric to $k$-dimensional Euclidean space, hyperbolic space with constant negative sectional curvature, a round real projective space, or a round sphere.
Proof.
Since $Y$ is $k$-dimensional, Theorem 2.18 implies that $\dim(\operatorname{Isom}(Y))\leq k(k+1)/2$. Since $\dim(\operatorname{Isom}(X,\mathcal{F}))$ is maximal, it follows from the proof of Theorem B (ii) that the subgroup $\Psi(\operatorname{Isom}(X,\mathcal{F}))\subset\operatorname{Isom}(Y)$ has dimension $k(k+1)/2$. Thus, $\operatorname{Isom}(Y)$ has dimension $k(k+1)/2$ and the conclusion follows from Theorem 2.18. ∎
We now determine the topology of the leaves and show that the submetry is a topological fibration, finishing the proof of item (iii) in Theorem  B.
Lemma 4.2.
Let $\pi\colon X\to Y$ be a proper submetry from an Alexandrov space $X$ of dimension $n$ onto an Alexandrov space $Y$ of dimension $0\leqslant k\leqslant n$. If
$\dim(\operatorname{Isom}(X,\mathcal{F}))=\frac{k(k+1)}{2}+\frac{(n-k)(n-k+1)}{2},$
then there exists a Lie group $G$ acting by isometries on $X$ such that the submetry $\pi$ is the composition of the submetry $\pi_{G}\colon X\to X/G$ and a submetry $\pi_{D}\colon X/G\to Y$ with discrete fibers. Moreover, all $G$-orbits are homeomorphic, and each leaf of $\pi$ is homeomorphic to a finite union of orbits of the $G$-action.
Proof.
Consider the group Lie group homomorphism $\Psi\colon\operatorname{Isom}(X,\mathcal{F})\to\operatorname{Isom}(Y)$ given by $\Psi(f)=\pi\circ f$ and set $G=\operatorname{Ker}(\Psi)$. As stated in the proof of Theorem B (ii), we may consider $\bar{x}\in X$ such that $\Sigma_{\bar{x}}(X)$ is isometric to the unit round $S^{n-1}$ and $\Sigma_{\pi(\bar{x})}(Y)$ is isometric to a unit round $S^{k-1}$.
From our hypotheses and the proof of Theorem B (ii), we observe that the orbit $G(\bar{x})$ has Hausdorff dimenstion equal to $n-k$, which is the Hausdorff dimension of the leaf $\pi^{-1}(\pi(\bar{x}))$.
We consider the space $X/G$ endowed with the orbit projection distance $d^{\ast}$. Note that we have a submetry $\pi_{G}\colon(X,d_{X})\to(X/G,d^{\ast})$. Furthermore, since the orbits of $G$ are contained in the leaves of $\pi$ (i.e., $G(x)\subset\pi^{-1}(\pi(x))$ for any $x\in X$), we obtain a well defined submetry $\pi_{D}\colon(X/G,d^{\ast})\to(Y,d_{Y})$ such that $\pi=\pi_{D}\circ\pi_{G}$. Note also that the Hausdorff dimension of the orbits $G(x)$ is bounded by the Hausdorff dimension of the leaves, which is bounded by $n-k$. Moreover, the leaf $L_{\bar{x}}=\pi^{-1}(\pi(\bar{x}))$ has Hausdorff dimension equal to $n-k$. From our hypotheses and the proof of Theorem B (ii), the dimension of the orbit $G(\bar{x})$ must also be $n-k$. Hence, the orbit space $X/G$ has Hausdorff dimension $n-(n-k)=k$. This implies that the submetry $\pi_{D}$ has discrete fibers, and thus $\pi_{D}^{-1}(\pi(x))=\sqcup_{i\in\mathbb{N}}\{x^{\ast}_{i}\}$. Since $\pi=\pi_{D}\circ\pi_{G}$, it follows that
$L_{\bar{x}}=\bigsqcup_{i\in\mathbb{N}}G(x_{i}).$
By [29, Theorem 7.2], the connected components of each fiber of $\pi$ have positive distance to one another.
Note now that, by Proposition 4.1, the space $Y$ is a Riemannian manifold. Thus, every point $y\in Y$ is a regular point of both $\pi$ and $\pi_{G}$. From the definition of regular point we conclude that given $x^{\ast}\in X/G$, the set of horizontal directions $H_{x^{\ast}}(\pi_{D})$ of $\pi_{D}$ is isometric to $\Sigma_{\pi_{G}(x^{\ast})}Y=S^{k-1}$. Moreover, since the fibers of $\pi_{D}$ is discrete, $H_{x^{\ast}}(\pi_{D})=\Sigma_{x^{\ast}}(X/G)$. This implies that all leaves of $\pi$ and $\pi_{G}$ are regular leaves. By Corollary 2.17, each leaf of $\pi$ is homeomorphic to $L_{\bar{x}}$, and each $G$-orbit is homeomomrphic to $G(\bar{x})$.
Let $U\subset G$ be an open neighborhood of the identity element $e\in G$ and set $U_{i}=\{gx_{i}\mid g\in U\}$, $i\in\mathbb{N}$. Then we have that $\{gU_{i}\mid g\in G,\>i\in\mathbb{N}\}$ is an open cover of $L_{\bar{x}}$. Moreover, by construction, $gU_{i}\cap hU_{j}=\varnothing$ for all $g,h\in G$ and $i\neq j$. Since $L_{\bar{x}}$ is compact, there exists a finite subcover of $\{gU_{i}\mid g\in G,\>i\in\mathbb{N}\}$. It follows then that $L_{\bar{x}}$ must be the finite union of disjoint orbits $G(x_{1}),\cdots,G(x_{N})$. ∎
Proposition 4.3.
Let $\pi\colon X\to Y$ be a proper submetry from an Alexandrov space $X$ of dimension $n$ onto an Alexandrov space $Y$ of dimension $0\leqslant k\leqslant n$. If
$\dim(\operatorname{Isom}(X,\mathcal{F}))=\frac{k(k+1)}{2}+\frac{(n-k)(n-k+1)}{2},$
then $(X,\mathcal{F})$ is foliated-homeomorphic to a fiber bundle over $Y$ whose fibers are isometric to a finite disjoint union of round spheres or a finite disjoint union of round projective spaces.
Proof.
Let $L$ be a regular leaf of the submetry $\pi\colon X\to Y$. Since $X$ has dimension $n$ and $Y$ is $k$-dimensional, $L$ has dimension $n-k$.
We start by proving that, under our hypotheses, the connected components of each leaf of the submetry are isometric to round spheres or round projective spaces.
As in the proof of Lemma 4.2, consider the group Lie group homomorphism $\Psi\colon\operatorname{Isom}(X,\mathcal{F})\to\operatorname{Isom}(Y)$ given by $\Psi(f)=\pi\circ f$, set $G=\operatorname{Ker}(\Psi)$, and let $\bar{x}\in X$ such that $\Sigma_{\bar{x}}(X)$ is isometric to the unit round $S^{n-1}$ and $\Sigma_{\pi(\bar{x})}(Y)$ is isometric to a unit round $S^{k-1}$. Each connected component $G(\bar{x})$ of $L_{\bar{x}}$ equipped with the intrinsic metric is a homogeneous space, locally compact, locally contractible, and since it is compact, each $1$-parameter subgroup of $G$ is rectifiable. Then, by [6, Theorem 7 I], $G(\bar{x})$ is a Finsler manifold. Moreover, since the spaces of directions of $L_{\bar{x}}$ are round $S^{n-k}$, the Minkowski cone of $G(\bar{x})$ is the Euclidean space $\mathbb{R}^{n-k}$. This implies that $G(\bar{x})$ with the intrinsic metric is a homogeneous Riemannian manifold. Therefore, $L_{\bar{x}}$ is a union of Riemannian manifolds and, hence, a compact Riemannian manifold with finitely many connected components. Moreover, by hypothesis, the isometry group of $L_{\bar{x}}$ has maximal dimension $(n-k)(n-k+1)/2$. By [28, Section II, Theorem 3.1], each connected component of $L_{\bar{x}}$ is isometric to a round $S^{n-k}$ or a round $\mathbb{R}P^{n-k}$. By Lemma 4.2, all connected components of $L_{\bar{x}}$ are homoeomorphic. Therefore, $L_{\bar{x}}$ is a finite union of either round spheres or round projective spaces.
Let us now show that $X$ is a topological manifold. Since every leaf of $\pi$ is a finite disjoint union of smooth manifolds and every point $x\in X$ is regular, we conclude that for any $x\in X$, the space of directions $\Sigma_{x}(X)$ is isometric to $V_{x}\ast H_{x}$. Recall that $V_{x}$ is the space of directions of $L_{x}$, and thus it is isomorphic to a unit round sphere $S^{n-k-1}$. The space $H_{x}$ is isometric to the space of directions $\Sigma_{\pi(x)}(Y)$, which is isometric to a unit round $S^{k-1}$. Therefore, the space of directions $\Sigma_{x}(X)$ is isometric to $S^{k-1}\ast S^{n-k-1}=S^{n-1}$. Hence, $X$ is a topological manifold.
We now show that $\pi\colon X\to Y$ is a fiber bundle. Note that, by [4], the map $\pi\colon X\to Y$ is in general at most a $C^{1,1}$ Riemannian submersion.
Recall taht $Y$ is connected. Fix $y_{0}\in Y$. Since $Y$ is a complete homogeneous Riemannian manifold, the injectivity radius is constant on $Y$. Fix $\varepsilon>0$ smaller than the injectivity radius, and consider the open cover $\{U_{\alpha}\}_{\alpha\in I}$ of $Y$ by normal balls of radius $\varepsilon$ centered at points $y_{\alpha}\in Y$. When $Y$ is a round sphere $S^{k}$ or a round projective space $\mathbb{R}P^{k}$, we avoid choosing $y_{\alpha}$ to be the conjugate point to $y_{0}$. For each $\alpha\in I$, let $c_{\alpha}\colon[0,1]\to Y$ be the unique minimizing geodesic joining $y_{0}$ to $y_{\alpha}$, set $F=\pi^{-1}(y_{0})\subset X$, and let $H=\operatorname{\mathrm{Homeo}}(F,F)=\{h\colon F\to F\mid\text{$h$ is a homeomorphism}\}$.
For $y\in Y$, set $F_{y}=\pi^{-1}(y)\subset X$ and let
$G_{y}=\operatorname{\mathrm{Homeo}}(F,F_{y})=\{h\colon F\to F_{y}\mid\text{$h$is a homeomorphism}\}.$
Let $P=\sqcup_{y\in Y}G_{y}$ and consider the map $\Upsilon\colon P\to Y$ given by $\Upsilon(g)=y$ if $g\in G_{y}$. We define an action of $H$ on $P$ as follows: given $h\in H$ and $g\in G_{y}\subset P$, we set $h\cdot g=g\circ h^{-1}\colon F\to F_{y}$, which is an element of $G_{y}$.
For each $y\in U_{\alpha}$, there is a unique minimizing geodesic $c_{y,\alpha}\colon[0,1]\to U_{\alpha}$ joining $y_{\alpha}$ to $y$. This geodesic is completely determined by its initial tangent vector $c_{y,\alpha}^{\prime}(0)\in T_{y_{\alpha}}Y$. Moreover, since $C(\Sigma_{x}(X))=C(H_{x})\times C(V_{x})$, for each unique minimizing geodesic $c\colon I\to Y$ in $Y$, given $x\in\pi^{-1}(c(0))$, there is a unique geodesic lift $\tilde{c}_{x}\colon I\to X$ by Lemma 2.14. Thus, we have a holonomy homeomorphism $h_{c}\colon\pi^{-1}(c(0))\to\pi^{-1}(c(1))$ given by $h_{c}(x)=\tilde{c}_{x}(1)$.
Define $s_{\alpha}\colon U_{\alpha}\to\pi^{-1}(U_{\alpha})$ by setting
$s_{\alpha}(y)=h_{c_{y,\alpha}}\circ h_{c_{\alpha}}\colon F\to F_{y}.$
Observe that this is a local section for the projection $\pi\colon P\to Y$ by construction; i.e., $\pi(s_{\alpha}(y))=y$ for all $y\in Y$.
Let us now show that the map $\Phi_{\alpha}\colon U_{\alpha}\times H\to\pi^{-1}(U_{\alpha})$ defined as
$\Phi_{\alpha}(y,h)=s_{\alpha}(y)\circ h^{-1}\colon F\to F_{y}$
is a bijection. Suppose that $\Phi_{\alpha}(y,h)=\Phi_{\alpha}(y^{\prime},h^{\prime})$. Then $y=\Upsilon(\Phi_{\alpha}(y,h))=\Upsilon(\Phi_{\alpha}(y^{\prime},h^{\prime}))=y^{\prime}$. Thus, $h^{-1}=s_{\alpha}(y)^{-1}\circ s_{\alpha}(y)\circ(h^{\prime})^{-1}=(h^{\prime})^{-1}$, which implies that $\Phi_{\alpha}$ is injective. Given a homeomorphism $g\colon F\to F_{y}$, consider $s_{\alpha}(y)\colon F\to F_{y}$. Then $h=g^{-1}\circ s_{\alpha}(y)$ is an element in $G$. Thus, $\Phi_{\alpha}(y,h)=s_{\alpha}(y)\circ h=g$, which implies that $\Phi_{\alpha}$ is onto. Thus, we may induce a topology on $\Upsilon^{-1}(U_{\alpha})$ making $\Phi_{\alpha}$ into a homeomorphism. Moreover, we have computed the inverse of $\Phi_{\alpha}$, which is given by
$\Phi_{\alpha}^{-1}(g)=(\Upsilon(g),g^{-1}\circ s_{\alpha}(\Upsilon(h))).$
For $U_{\alpha}\cap U_{\beta}\neq\emptyset$ we have:
$\Phi_{\beta}^{-1}\circ\Phi_{\alpha}(y,h)=\Phi_{\beta}^{-1}(s_{\alpha}(y)\circ h^{-1}))=(y,h\circ s_{\alpha}(y)^{-1}\circ s_{\beta}(y)).$
Thus, transition functions are of the form $t_{\alpha\beta}=s_{\alpha}^{-1}\circ s_{\beta}\colon U_{\alpha}\cap U_{\beta}\to H$ and are continuous by construction, and satisfy the cocycle condition.
Therefore, $\Upsilon\colon P\to Y$ is a continuous principal bundle. We see that $X$ is the associated fiber bundle via the homeomorphism $\varphi\colon P\times_{H}F\to X$ given by $\varphi([g,v])=g(v)$. ∎